Consider an exchange traded put option to sell 200 shares for $60 per share (strike price) for company ABC. Suppose the company ABC announces a 3 for 2 stock split, please answer the following questions.
(a) What is the new strike price after the stock split ?
(b) What is the number of shares that can be sold after the stock split ?
Answer #1

For every 2 shares, the investor gets three shares.

The initial notional value of the contract = 200 * $60 = $12,000

(a) The new strike price after the stock split = Old strike price / split ratio

The new strike price after the stock split = 60/3 = $20

(b) Number of shares that can be sold after the stock split = Initial number of shares * split ratio

Number of shares that can be sold after the stock split = 200 * 3 = 600

The notional value after the split = 600 * $20 = $12,000

Can you please upvote? Thank you :-)